# SMITE: Segment Me In TimE

### v1

@@ Abstract
Amirhossein Alimohammadi1, Sauradip Nag1, Saeid Asgari Taghanaki1,2,
Andrea Tagliasacchi1,3,4, Ghassan Hamarneh1, Ali Mahdavi Amiri1
1Simon Fraser University 2Autodesk Research 3University of Toronto 4Google DeepMind
@@ 2 Related Work
Video segmentation. Video segmentation methods can be categorized as video semantic segmentation (VSS) (Zhu et al., 2024; Zhang et al., 2023a; b; Li et al., 2024; Ke et al., 2023; Wang et al., 2024), video instance segmentation (VIS) (Yang et al., 2019) and video object segmentation (VOS) (Xie et al., 2021; Wang et al., 2021b; Cheng et al., 2021a; b; Bekuzarov et al., 2023). VSS and VIS extends image segmentation to videos, assigning pixel labels across frames while maintaining temporal consistency despite object deformations and camera motion. VOS, in contrast, focuses on tracking and isolating specific objects throughout the video. Both tasks leverage temporal correlations through techniques like temporal attention (Mao et al., 2021; Wang et al., 2021a), optical flow (Xie et al., 2021; Zhu et al., 2017), and spatio-temporal memory (Wang et al., 2021b; Cheng & Schwing, 2022). Recent efforts, such as UniVS (Li et al., 2024), proposes unified models for various segmentation tasks, utilizing prior frame features as visual prompts. However, these methods struggle with fine-grained part segmentation and generalization to unseen datasets (Zhang et al., 2023b). Bekuzarov et al. (2023) leverage a spatio-temporal memory module and a frame selection mechanism to achieve high-quality video part segmentation with partial annotations. However, it requires frame annotations from the same video complicating video segmentation at scale. In contrast, we only need segmentation references for a few arbitrary selected images and we can segment an unseen given video respecting the segementation references. Therefore, per video manual annotation is not needed in our method.
@@ 3 Preliminaries
Latent diffusion models and WAS maps. Latent Diffusion Models perform the denoising operation on the latent space of a pretrained image autoencoder. Each latent pixel corresponds to a patch in the generated image. Starting from pure random noise $z_{T}$, at each timestep $t$, the current noisy latent $z_{t}$ is passed through a denoising UNet $\epsilon_{\theta}$, which is trained to predict the current noise $\epsilon_{\theta}(z_{t},y,t)$ using text prompt $y$. In each block, the UNet employs residual convolution layers to generate intermediate features, which are then fed into attention layers. These attention layers average various values based on pixel-specific weights. The cross-attention layers (denoted by $A_{ca}$) incorporate semantic contexts from the prompt encoding, whereas the self-attention layers (denoted by $A_{sa}$) leverage global information from the latent representation itself. SLiMe (Khani et al., 2024) demonstrated that text embeddings can be learned from a few images to segment other unseen images by leveraging both cross attention and self-attention layers in a novel representation called Weighted Accumulated Self-Attention (WAS) map, $S_{\text{WAS}}$. This is defined as follows:
@@ 4.1 Problem Setting
Our framework. We use Stable Diffusion’s (SD) semantic knowledge to learn the segments defined by few images and then generalize them to the video of the subject in any pose, color or size. This implies that the model needs to share information across frames to enforce temporal consistency. Differently from the UNet structure used in Khani et al. (2024), we apply an inflation to the T2I models across the temporal dimension (Wu et al., 2022) to enable temporal attention across all video frames. Also, we incorporate a tracking module combined with low-frequency regularization to enhance spatio-temporal consistency across frames. The overall inference pipeline of our model, SMITE, is illustrated in Fig. 2.
@@ 4.2 Learning generalizable segments
Network fine-tuning. Only learning text embeddings fails to capture complex granularities. To extract more customizable fine-grained segments from the video, we need to fine-tune the existing SD weights (denoted by $\theta$) using the available provided segment annotations. As shown in the inset figure, optimizing solely the text embeddings struggles with asymmetrical segmentation (e.g., segmenting only one eye). We hypothesize that such issues arises because the model may get stuck in local minima when relying exclusively on text embeddings for optimization. To mitigate this, we update the cross-attention layers $A_{ca}$ in the UNet along with the text-embedding but do so in two phases. First, the model is frozen, while the text embeddings corresponding to the segments denoted by $\mathcal{T}$ are optimized using a high learning rate. Thus, an initial embedding ($\mathcal{T}^{*}$) is achieved quickly without detracting from the generality of the model, which then serves as a good starting point for the next phase. Second, we unfreeze the cross-attention weights and optimize them along with the text embeddings, using a significantly lower learning rate. This gentle fine-tuning of the cross-attention and the embeddings enables faithful generation of segmentation masks with varied granularity. We use the same combination of losses ($\mathcal{L}_{\text{CE}}$, $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\text{LDM}}$) used in SLiMe (Khani et al., 2024) for network finetuning. This results in an optimized SMITE model (denoted by $\Psi_{\theta}^{*}$) which can generate segments that is generalizable across different videos.
@@ 4.3 Temporal Consistency
To enhance temporal consistency, we use temporal attention in SMITE’s Inflated UNet denoiser. While it improves temporal consistency compared to independent frame processing (Fig. 3(a)), inconsistencies remain due to the need for segmenting at different granularities, often with imprecise boundaries. These inconsistencies can cause flickering or unnatural transitions (Fig. 3(b)).
@@
Figure 3: Best viewed in Adobe Acrobat.
@@
Segment tracking and voting. The first step to ensure segment consistency is tracking the segments across time. Point tracking methods like CoTracker (Karaev et al., 2023) are well-suited for our approach because they use point correspondences to minimize pixel drift over time. However, since our segments come from attention maps, tracking needs to occur directly on these maps. Since CoTracker is trained on spatial domains, we first apply it on the frames of video $\mathcal{V}$ using a sliding window, project the results onto the attention maps via simple scaling, and then update the attention maps accordingly. A pixel’s label in the attention map is updated based on the most frequent label it receives across the visible frames in the window. Formally, for a pixel with coordinates $(x_{t},y_{t})$ on the frame latent at timestep $t$, its coordinates on all subsequent frame latents in the window can be derived from the tracker. The coordinates are linked, and the trajectory sequence can be presented as:
@@
where $\mathcal{P}$ is the tracker, $\phi(.)$ is the scaling operator, and $w$ denotes the length of the sliding window. We choose the point corresponding to the center frame of the window $(x_{t},y_{t})$ as the first point and track bidirectional correspondences within the window. Each of these trajectory points $(x,y)$ is assigned a segment label $\hat{Y}$. However, certain pixels in the window may disappear and reappear over time as shown in Fig 4. To handle such scenarios, we discard labels for which the tracking pixel is invisible and then use temporal voting to update its segment labels. Therefore, we update the WAS maps (denoted by $\mathcal{S}_{\text{WAS}}$) corresponding to the tracking pixel $(x_{t},y_{t})$ and use the following formula to assign the labels:
@@
where $F$ denotes the most frequent label in the window. After correcting the unstable pixel tracking over multiple-step denoising, we obtain a consistent set of segments ($\mathcal{S}_{tracked}$). Note that the window size $w$ is small (e.g, seven), hence to track the segments for the entire video, we need to slide the windows and track in an iterative fashion. More details are provided in the supplementary material.
@@
Low-pass regularization. When tracking is applied, label modifications should be repeated through several denoising steps to ensure full temporal label propagation and achieve better consistency. This process may cause deviations (Fig. 3) from the original segmentation provided by the reference images that is captured in the segments ($\mathcal{S}$) in the initial denoising step. Our goal is to preserve the overall structure of the segments $\mathcal{S}$ from the WAS maps while smoothing boundary transitions for temporal consistency via tracking. Low Pass Filters (LPF) have been effective in video generation to maintain temporal correlations and spatial frame structure (Wu et al., 2023; Si et al., 2024). Therefore, we use an LPF on the finally denoised segments $\mathcal{S}$ ensuring the structure of the LPF on the segments $\mathcal{S}$ predicted in the initial denoising step (denoted by $\mathcal{S}_{ref}$) is respected by the following function:
@@
where $\omega(S)=\mathcal{H}_{l}\odot\text{DCT}_{3D}(\text{S})$, $\text{DCT}_{3D}$ refers to Discrete Cosine Transform and $\mathcal{H}_{l}$ refers to an LPF. This ensures that the segments structure is progressively corrected across the video frames at the end of the denoising phase as illustrated in Fig. 3.
@@ 4.4 Spatio-Temporal Guidance
Our overall energy function is minimized by backpropagating through the diffusion process, as described in Chen et al. (2023a); Safaee et al. (2023), which updates the latent representation to achieve more consistent segmentation over time:
@@
where $\lambda_{\text{Tracking}}$ and $\lambda_{\text{Reg}}$ are the coefficients for the tracking and reference loss functions, respectively. Optimizing this function encourages better spatial and temporal consistency in the WAS attention maps (S). Specifically, we update the latent $z_{t}$ using gradient descent on $\mathcal{E}_{\text{Total}}$:
@@ 5 Results and Experiments
Quantitative comparison. Since there are no available few-shot video part segmentation methods that can be applied to our setting, we build the following two state-of-the-art baselines for quantitative comparison: (1) Few-shot Image segmentation based approach, in which SLiMe (Khani et al., 2024) is applied to the video frame-by-frame. We term this as Baseline-I. (2) Grounded SAM2 (Ren et al., 2024; Ravi et al., 2024), a recently introduced zero-shot foundation model for video segmentation approach which is applied directly on the video. For Baseline-I, we applied the same training procedure and used SMITE-50 dataset. As apparent in Tab. 1, our method produces the most accurate results in comparison with other methods across all categories and all metrics.
@@
Faces Horses Cars Non-Text
Methods $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU $\textbf{F}_{meas.}$ mIOU
Baseline-I 0.81 72.95 0.64 65.48 0.57 61.38 0.67 66.69
GSAM2 0.73 63.28 0.76 72.76 0.64 63.56 - -
Ours 0.89 77.28 0.79 75.09 0.82 75.10 0.77 73.08
@@
Qualitative comparisons. Qualitative comparisons are presented in Fig. 6. GSAM2 struggles to locate the boundaries accurately and produces coarse segments. SLiMe can usually segment the first frame accurately but struggles to preserve temporal consistency. Our method produces the best segmentation maps in terms of segmentation quality and temporal consistency. The maps have sharper boundaries and clean clusters compared to other methods. Our SMITE performs better than its SLiMe counterpart. In Fig. 7, we present results from other categories, demonstrating the versatility of our model in handling various object categories. As shown, despite significant differences in pose, expression, gender, and other properties between the video and the annotated images, our method still delivers high-quality results. This is particularly evident in the pineapple example, where the object is cut in half, yet the method successfully tracks it and produces accurate segmentation. For more challenging cases, we include results from additional difficult videos in Fig. 8. In one instance, the ice cream cone is occluded by a paper napkin, and the ice cream itself is obscured and blended by a face, yet SMITE is still able to generate correct results. Furthermore, the turtle nearly blends with the background in terms of color and visual patterns, but our method successfully tracks the segments.
Ablation study. To demonstrate the effectiveness of our design choices, we conducted three ablation studies. The first ablation on the Car dataset (Tab. 3), illustrates that combining $\mathcal{E}_{\text{Tracking}}$ and $\mathcal{E}_{\text{Reg}}$ improves the results. It’s important to note that tracking without low pass regularization leads to unsatisfactory outcomes, as results tend to align too closely with the voting method. Tab. 3 shows the impact of cross attention tuning also on the Car dataset. By optimizing both cross attention and text embedding, we achieve the highest accuracy, demonstrating the effectiveness of our tuning strategy. We have also ablated the number of images for training our few shot setting $k=1/5/10$. While our performance improves with more training images, our one shot setting still performs reasonably well (Tab. 5).
Table 2: Loss ablation. When two losses are combined, the best performance is achieved.
$\mathcal{E}_{\text{Tracking}}$ $\mathcal{E}_{\text{Reg}}$ mIOU
69.85
75.10
Table 3: Cross attention optimization improves the quality of segmentation.
Finetuning Strategy mIOU
Text Embedding 74.00
Cross Attention + Text 75.10
@@
User study. We conduct a user study because human judgment is best for assessing perception when both temporal consistency and faithfulness to segmentations are important. We collected 16 segmented videos (4 from Non-Text category) and asked 40 participants to rank the methods (i.e., 1 best and 3 worst) based on segmentation quality (i.e., fidelity to the segmentation reference) and motion consistency (i.e., reduced flicker), encouraging them to prioritize segmentation quality in their evaluations in terms of a tie in motion consistency. We do not report the scores of GSAM2 on Non-Text segments as the segments in the videos are not describable by texts. Tab. 5 demonstrates that SMITE achieves higher preference both for textual and non-textual segmentation.
@@
Table 4: Few shot ablation on cars in SMITE-50. The performance increases with more training images but still performs well in one shot setting.
@@
1-shot 63.03
5-shot 71.55
10-shot 75.10
@@
Baseline-1 2.58 2.37
GSAM2 2.13 -
Ours 1.19 1.10
6 Conclusion
In this work, we introduce SMITE, a video segmentation technique that supports flexible granularity segmentation of objects within a video. SMITE leverages the semantic knowledge of a pre-trained text-to-image diffusion model to segment a video with minimal additional training (i.e., only a few images), while utilizing the temporal consistency of the video to maintain motion consistency within the segments. Notably, SMITE can be trained with a few images that are not necessarily from the video, yet it effectively segments objects within an unseen video. To better show the capabilities of our method, we also collected a flexible granularity dataset called SMITE-50 that will be publicly available along with our code. Through various quantitative and qualitative experiments, as well as user studies, we demonstrate the effectiveness of our method and its components. However, our method still faces limitations that motivate future work. Specifically, SMITE does not perform well when the target objects or segments are too small, and its performance degrades when dealing with with low video resolution. Addressing these issues presents interesting avenues for future research. Additionally, while we use Co-tracker to track segments, exploring other tracking systems and evaluating their performance would be valuable.

### v2

@@ 2 Related Work
Video segmentation. Video segmentation methods can be categorized as video semantic segmentation (VSS) (Zhu et al., 2024; Zhang et al., 2023a; b; Li et al., 2024; Ke et al., 2023; Wang et al., 2024), video instance segmentation (VIS) (Yang et al., 2019) and video object segmentation (VOS) (Xie et al., 2021a; Wang et al., 2021b; Cheng et al., 2021c; d; Bekuzarov et al., 2023). VSS and VIS extends image segmentation to videos, assigning pixel labels across frames while maintaining temporal consistency despite object deformations and camera motion. VOS, in contrast, focuses on tracking and isolating specific objects throughout the video. Both tasks leverage temporal correlations through techniques like temporal attention (Mao et al., 2021a; Wang et al., 2021a), optical flow (Xie et al., 2021a; Zhu et al., 2017), and spatio-temporal memory (Wang et al., 2021b; Cheng & Schwing, 2022). Recent efforts, such as UniVS (Li et al., 2024), proposes unified models for various segmentation tasks, utilizing prior frame features as visual prompts. However, these methods struggle with fine-grained part segmentation and generalization to unseen datasets (Zhang et al., 2023b). Bekuzarov et al. (2023) leverage a spatio-temporal memory module and a frame selection mechanism to achieve high-quality video part segmentation with partial annotations. However, it requires frame annotations from the same video complicating video segmentation at scale. In contrast, we only need segmentation references for a few arbitrary selected images and we can segment an unseen given video respecting the segementation references. Therefore, per video manual annotation is not needed in our method.
@@ 3 Preliminaries
Latent diffusion models and WAS maps. Latent Diffusion Models perform the denoising operation on the latent space of a pretrained image autoencoder. Each latent pixel corresponds to a patch in the generated image. Starting from pure random noise $z_{T}$, at each timestep $t$, the current noisy latent $z_{t}$ is passed through a denoising UNet $\left(\epsilon_{\theta}\right)$ , which is trained to predict the current noise $\epsilon_{\theta}(z_{t},y,t)$ using text prompt $y$. In each block, the UNet employs residual convolution layers to generate intermediate features, which are then fed into attention layers. These attention layers average various values based on pixel-specific weights. The cross-attention layers (denoted by $A_{ca}$) incorporate semantic contexts from the prompt encoding, whereas the self-attention layers (denoted by $A_{sa}$) leverage global information from the latent representation itself. SLiMe (Khani et al., 2024) demonstrated that text embeddings can be learned from a few images to segment other unseen images by leveraging both cross attention and self-attention layers. In fact, each segment of an image is linked with a text token that is optimized to align its associated cross-attention and self-attention maps with the corresponding segment in the provided training image. For segmentations with multiple levels of granularity, multiple tokens are optimized, with each token representing a distinct segment. The optimization process of attention maps is captured in a novel representation called Weighted Accumulated Self-Attention (WAS) map, $S_{\text{WAS}}$. This is defined as follows:
@@ 4.1 Problem Setting
Our framework. We use Stable Diffusion’s (SD) semantic knowledge to learn the segments defined by few images and then generalize them to the video of the subject in any pose, color or size. This implies that the model needs to share information across frames to enforce temporal consistency. Differently from the UNet structure used in Khani et al. (2024), we apply an inflation to the T2I models across the temporal dimension (Wu et al., 2022) to enable temporal attention across all video frames. Also, we incorporate a tracking module combined with low-frequency regularization to enhance spatio-temporal consistency across frames. The overall inference pipeline of our model, SMITE, is illustrated in Fig 2.
@@ 4.2 Learning generalizable segments
Network fine-tuning. Only learning text embeddings fails to capture complex granularities. To extract more customizable fine-grained segments from the video, we need to fine-tune the existing SD weights (denoted by $\theta$) using the available provided segment annotations. As shown in the inset figure, optimizing solely the text embeddings struggles with asymmetrical segmentation (e.g., segmenting only one eye). We hypothesize that such issues arises because the model may get stuck in local minima when relying exclusively on text embeddings for optimization. To mitigate this, we update the cross-attention layers $A_{ca}$ in the UNet along with the text-embedding but do so in two phases. First, the model is frozen, while the text embeddings corresponding to the segments denoted by $\mathcal{T}$ are optimized using a high learning rate. Thus, an initial embedding ($\mathcal{T}^{*}$) is achieved quickly without detracting from the generality of the model, which then serves as a good starting point for the next phase. Second, we unfreeze the cross-attention weights and optimize them along with the text embeddings, using a significantly lower learning rate. This gentle fine-tuning of the cross-attention and the embeddings enables faithful generation of segmentation masks with varied granularity. For both training phases, we use the same combination of losses ($\mathcal{L}_{\text{CE}}$, $\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}_{\text{LDM}}$) used in SLiMe (Khani et al., 2024). This results in an optimized SMITE model (denoted by $\Psi_{\theta}^{*}$) generating generalizable segment across different videos.
@@ 4.3 Temporal Consistency
To enhance temporal consistency, we use temporal attention in SMITE’s Inflated UNet denoiser. While it improves temporal consistency compared to independent frame processing (Fig 3(a)), inconsistencies remain due to the need for segmenting at different granularities, often with imprecise boundaries. These inconsistencies can cause flickering or unnatural transitions (Fig 3(b)).
@@
Figure 3: Video best viewed in Acrobat.
@@
Segment tracking and voting. The first step to ensure segment consistency is tracking the segments across time. Point tracking methods like CoTracker (Karaev et al., 2023) are well-suited for our approach because they use point correspondences to minimize pixel drift over time. However, since our segments come from attention maps, tracking needs to occur directly on these maps. Since CoTracker is trained on spatial domains, we first apply CoTracker (denoted by $\mathcal{P}$) on the frames of video $\mathcal{V}$ bidirectionally as $\{X,Y\}=\mathcal{P}(\mathcal{V})$ where $X,Y$ are set of tracked pixel coordinates. To project these tracking onto the the attention maps, we use a sclaing operator $\phi(.)$ and linearly scale the pixel trajectories onto the attention space. We process the entire video in a temporal sliding window fashion having window length $w$. For each of such temporal window, we then leverage the projected trajectories and update the attention pixel’s label based on the most frequent label it receives across the visible frames in the window. Formally, for a pixel with coordinates $(x_{t},y_{t})$ on the attention map latent at timestep $t$, its coordinates on all subsequent latents in the window can be derived from the tracker. The coordinates are linked, and the trajectory sequence can be presented as:
@@
We choose the point corresponding to the center frame of the window $(x_{t},y_{t})$ as the first point and leverage bidirectional correspondences within the window. Each of these trajectory points $(x,y)$ is assigned a segment label $\hat{Y}$. However, certain pixels in the window may disappear and reappear over time as shown in Fig 4. To handle such scenarios, we discard labels for which the tracking pixel is invisible and then use temporal voting to update its segment labels. Therefore, we update the WAS maps (denoted by $\mathcal{S}_{\text{WAS}}$) corresponding to the tracking pixel $(x_{t},y_{t})$ and use the following formula to assign the labels:
@@
where $F$ denotes the most frequent label in the window.
After correcting the unstable pixel tracking over multiple-step denoising, we obtain a consistent set of segments ($\mathcal{S}_{tracked}$). Note that the window size $w$ is small (e.g, seven), hence to track the segments for the entire video, we need to slide the windows and track in an iterative fashion. More details are provided in the supplementary material.
@@
Low-pass regularization. When tracking is applied, label modifications should be repeated through several denoising steps to ensure full temporal label propagation and achieve better consistency. This process may cause deviations (Fig 3) from the original segmentation provided by the reference images that is captured in the segments ($\mathcal{S}$) in the initial denoising step. Our goal is to preserve the overall structure of the segments $\mathcal{S}$ from the WAS maps while smoothing boundary transitions for temporal consistency via tracking (refer to Fig 8). Low Pass Filters (LPF) have been effective in video generation to maintain temporal correlations and spatial frame structure (Wu et al., 2023; Si et al., 2024). Therefore, we use an LPF on the finally denoised segments $\mathcal{S}$ ensuring the structure of the LPF on the segments $\mathcal{S}$ predicted in the initial denoising step (denoted by $\mathcal{S}_{ref}$) is respected by the following function:
@@
where $\omega(S)=\mathcal{H}_{l}\odot\text{DCT}_{3D}(\text{S})$, $\text{DCT}_{3D}$ refers to Discrete Cosine Transform and $\mathcal{H}_{l}$ refers to an LPF. This ensures that the segments structure is progressively corrected across the video frames at the end of the denoising phase as illustrated in Fig 3.
@@ 4.4 Spatio-Temporal Guidance
Our overall energy function is minimized at test time by backpropagating through the diffusion process, as described in Chen et al. (2023a); Safaee et al. (2023), which updates the latent representation to achieve more consistent segmentation over time:
@@
where $\lambda_{\text{Tracking}}$ and $\lambda_{\text{Reg}}$ are the coefficients for the tracking and reference loss functions, respectively. We have two different energy functions: $\mathcal{E}_{\text{Tracking}}$ which reduces flicker and ensures segment consistency over time, while $\mathcal{E}_{\text{Reg}}$ which maintains correct semantics while preventing segment drift. Thus, optimizing $\mathcal{E}_{\text{Total}}$ encourages better spatial and temporal consistency in the WAS attention maps (S). Specifically, we update the latent $z_{t}$ using gradient descent on $\mathcal{E}_{\text{Total}}$:
@@ 5 Results and Experiments
Quantitative comparison. Since there are no available few-shot video part segmentation methods that can be applied to our setting, we build the multiple baselines for quantitative comparison. First, a few-shot image segmentation based approach, in which SLiMe (Khani et al., 2024) is applied to the video frame-by-frame. We term this as Baseline-I. Second, Grounded SAM2 (Ren et al., 2024; Ravi et al., 2024), a recently introduced zero-shot foundation model for video segmentation approach which is applied directly on the video. For Baseline-I, we applied the same training procedure and used SMITE-50 dataset. Another baseline, Baseline-II is to use SLiMe and CoTracker simultaneously. However, directly applying tracking in the pixel space does not produce improved results, as it can lead to drift in segmentation. In contrast, SMITE leverages tracking on the attention maps, focusing on where the object and its parts are located, which minimizes the impact of pixel color shifts. Additionally, we have incorporated a regularization mechanism into our tracking process to further mitigate such drifts, ensuring alignment with the true semantic segmentations. We have also provided another baseline, Baseline-III, which is the modification of SLiMe by only changing its UNet to the Inflated UNet. As shown in Tab. 1, our method produces the most accurate results compared to other baselines across all categories and metrics.
@@
Faces Horses Cars Non-Text Mean
Methods $\textbf{F}_{\text{meas.}}$ mIOU $\textbf{F}_{\text{meas.}}$ mIOU $\textbf{F}_{\text{meas.}}$ mIOU $\textbf{F}_{\text{meas.}}$ mIOU $\textbf{F}_{\text{meas.}}$ mIOU
Baseline-I 0.81 72.95 0.64 65.48 0.57 61.38 0.67 66.69 0.67 66.62
Baseline-II 0.78 71.08 0.62 64.01 0.65 61.95 0.65 64.96 0.67 65.50
Baseline-III 0.82 73.25 0.64 67.48 0.70 68.49 0.66 68.22 0.70 69.36
GSAM2 0.73 63.28 0.76 72.76 0.64 63.56 - - 0.71 66.53
Ours 0.89 77.28 0.79 75.09 0.82 75.10 0.77 73.08 0.82 75.14
@@
Qualitative comparisons. Qualitative comparisons are presented in Fig. 6. GSAM2 struggles to locate the boundaries accurately and produces coarse segments. SLiMe can usually segment the first frame accurately but struggles to preserve temporal consistency. Our method produces the best segmentation maps in terms of segmentation quality and temporal consistency. The maps have sharper boundaries and clean clusters compared to other methods. Our SMITE performs better than its SLiMe counterpart. In Fig. 7, we present results from other categories, demonstrating the versatility of our model in handling various object categories. As shown, despite significant differences in pose, expression, gender, and other properties between the video and the annotated images, our method still delivers high-quality results. This is particularly evident in the pineapple example, where the object is cut in half, yet the method successfully tracks it and produces accurate segmentation. More qualitative results are provided in Appendix 7.4.
Ablation study. To demonstrate the effectiveness of our design choices, we conducted a comprehensive ablation study showing the impact of each component of our method one by one on the entire SMITE-50 dataset. Tab. 2 shows that our final setting provides the highest accuracy while the impact of each component is noticeable. Note that the low-frequency regularization (LP Reg) helps maintain spatial structure and prevents segment drift when tracking is applied to $S_{WAS}$. Without tracking, LP Reg has minimal effect since no segment drift occurs in WAS maps (thus change of energy in $\mathcal{E}_{\text{Reg}}$ is almost zero), so we do not ablate LP Reg separately without tracking. We have also performed an ablation on the number of training images illustrated in Tab. 4. As expected the accuracy improves when the number of training samples increases.
Table 2: Network Ablation on SMITE-50 dataset. Ablations on the entire SMITE-50 dataset. We respectively show the impact of inflated UNet (IUNet), cross attention tuning (CA Tuning), Tracking, and Low-pass (LP) regularization.
Method Baseline I (SLiMe) +IUNet (Baseline III) +CA Tuning +Tracking +LP Reg (SMITE)
mIOU 66.22 69.36 71.70 72.64 75.14
@@
User study. We conduct a user study because human judgment is best for assessing perception when both temporal consistency and faithfulness to segmentations are important. We collected 16 segmented videos (4 from Non-Text category) and asked 25 participants to rank the methods (i.e., 1 best and 5 worst) based on segmentation quality (i.e., fidelity to the segmentation reference) and motion consistency (i.e., reduced flicker), encouraging them to prioritize segmentation quality in their evaluations in terms of a tie in motion consistency. We do not report the scores of GSAM2 on Non-Text segments as the segments in the videos are not describable by texts. Tab. 4 demonstrates that SMITE achieves higher preference both for textual and non-textual segmentation.
@@
Table 3: Few shot ablation on SMITE-50. The performance increases with more training images but still performs well in one shot setting.
@@
        1-shot         58.55
        5-shot         73.88
        10-shot         75.14
@@
Baseline-I 3.59 2.41
Baseline-II 3.42 3.24
Baseline-III 2.98 3.32
GSAM2 3.37 -
Ours 1.46 1.03
6 Conclusion, Limitations, Future Work
In this work, we introduce SMITE, a video segmentation technique that supports flexible granularity segmentation of objects within a video. SMITE leverages the semantic knowledge of a pre-trained text-to-image diffusion model to segment a video with minimal additional training (i.e., only a few images), while utilizing the temporal consistency of the video to maintain motion consistency within the segments. Notably, SMITE can be trained with a few images that are not necessarily from the video, yet it effectively segments objects within an unseen video. To better show the capabilities of our method, we also collected a flexible granularity dataset called SMITE-50 that will be publicly available along with our code. Through various quantitative and qualitative experiments, as well as user studies, we demonstrate the effectiveness of our method and its components. However, our method still faces limitations that motivate future work. Firstly, although, our regularization ensures that tracking does not dominate the segmentation process or deviate from the anticipated semantic segmentations, since SMITE utilizes an off-the-shelf tracking technique, its tracking mechanism and voting system are still limited by the performance of the underlying tracking system to some extent. Secondly, the generation of WAS maps involves the use of low-resolution cross-attention mechanisms, which may result in reduced accuracy when segmenting fine details or small segments, such as tiny strings or minor components. Thirdly, it cannot offer an interactive segmentation since it involves the diffusion and optimization processes. In this work, we did not particularly focus on optimizing training and inference time but this can be an interesting future work.